OpenStudy (anonymous):

Use the general formula to write the area A of the enclosure as a function of w (that is, only the variable w should appear in the function). it should be a quadratic function

OpenStudy (anonymous):

What enclosure?

myininaya (myininaya):

what is the enclosure? are we enclosing a rectangle or something? or can we enclose whatever we want?

OpenStudy (anonymous):

hold on im trying to get the file of the actuall question its part of a project

OpenStudy (anonymous):

Its to question number 5

OpenStudy (anonymous):

I am completely lost

myininaya (myininaya):

maybe this will help

OpenStudy (anonymous):

Okay, so you have 682 feet of fencing to make up two "width sides" and one "length side" this translates to the equation: 682 = 2W + L You know from basic geometry that the area of a rectangle is length times width: A = L*W What you need to do is solve the first equation for L in terms of W: 682 = 2W + L L = 682 -2W Then substitute that L value into the second function so that you have A in terms of just W: A = L*W A = (682 -2W)W

OpenStudy (anonymous):

anyone?

myininaya (myininaya):

?

myininaya (myininaya):

anyone what?

OpenStudy (anonymous):

the site froze on me im taking a look at it now

myininaya (myininaya):

ok lol

OpenStudy (anonymous):

now if i can get the window correct on my ti 84 :P

OpenStudy (anonymous):

any help with that?

myininaya (myininaya):

go to Y=

myininaya (myininaya):

type 682x-2x^2

myininaya (myininaya):

then go to window

OpenStudy (anonymous):

yeah i got the equation in its just getting a clear shot of the parabola

myininaya (myininaya):

go to where it says x min and type 100 go to where it says x max and type 200 go to where it says y min and type 80 go to where it says y max and type 60000 and see if it looks viewable

myininaya (myininaya):

try going to the highest point on the curve and approximate what it is

OpenStudy (anonymous):

it works but im going to need both sides of the parabola can i just extend the xmin

OpenStudy (anonymous):

holy crap I got it

myininaya (myininaya):

no you are fine but you can if you want to

OpenStudy (anonymous):

you are amazing!

myininaya (myininaya):

did you use the maximum feature on your calculator?

OpenStudy (anonymous):

ximn -100 xmax 400 ymin 80 ymax 60000

myininaya (myininaya):

no thats not what i meant

myininaya (myininaya):

do you see calc?

OpenStudy (anonymous):

i havent taken advantage of that yet

myininaya (myininaya):

lets do it

OpenStudy (anonymous):

i got the max

myininaya (myininaya):

do you see maximum

OpenStudy (anonymous):

yeah did the whole left and right bound

OpenStudy (anonymous):

thats going to give me the best area correct?

myininaya (myininaya):

oh nice i got (170.50001,58140.5) this is just an appoximation I arleady found the exact on that pdf file

myininaya (myininaya):

your approximation doesn't have to be mine but both of ours should be pretty close to the same thing is yours close to my?

OpenStudy (anonymous):

its exact

myininaya (myininaya):

k now you remember the file i gave you right? that vertex of the parabola i found by putting the parabola in vertex form gave me the exact (max width,max area)

OpenStudy (anonymous):

what file

myininaya (myininaya):

scroll up if you haven't seen the file and click on it

OpenStudy (anonymous):

looking at it now

myininaya (myininaya):

you remember when you uploaded a file i uploaded a file like a few minutes afterwards

OpenStudy (anonymous):

yeah i see it

OpenStudy (anonymous):

if i have any more questions ill post them

myininaya (myininaya):

ok make a new post though because i'm about to leave or do it so you can get someone's attention and bring them back here to help k?

myininaya (myininaya):

or 3wood might come back

myininaya (myininaya):

gn guys and peace

OpenStudy (anonymous):

thanks for all the help both of you are life savers